Find the equation for the line that passes the point(-1,-5) and that is parallel to the line with the equation y=1/3×+2
Answer by sachi(548)

y=1/3×+2=mx+c
so by comparision m=slope of the give line=1/3
now the new line is parallel to the above line
so the slope of new line=1/3 & it passes thru (-1,-5)
so the required eqn is [y-(-5)]/[x-(-1)]=1/3
or (y+5)/(x+1)=1/3
or 3(y+5)=x+1
or 3y-x+14=0